OpenStudy (anonymous):

solve for y, where y is a real number sqrt(-2y+24)=y

OpenStudy (anonymous):

Square both sides to get -2y+24=y^2

OpenStudy (anonymous):

yes thats what i did

OpenStudy (anonymous):

Then get everything on one side so y^2+2y-24=0

OpenStudy (anonymous):

ohh yeaa

OpenStudy (anonymous):

Then factor to get (y+6)(y-4)=0

OpenStudy (anonymous):

set (y+6) = 0 then (y-4)=0

OpenStudy (anonymous):

excellent

OpenStudy (anonymous):

y=-6,4

OpenStudy (anonymous):

sweet i remember haha thanks @AEYES13

OpenStudy (anonymous):

@ludwig457 Not a problem!

OpenStudy (anonymous):

btw its y=4! thanks again

OpenStudy (anonymous):

because -6 wont work lol

OpenStudy (anonymous):

Alright sweet as long as you got the right answer

OpenStudy (anonymous):

yessir